OpenStudy (anonymous):

Identify the excluded value for each function. 35. y=3/x +4

OpenStudy (anonymous):

\[y=\frac{ 3 }{ x+4 }\]

OpenStudy (anonymous):

x+4 notequal 0 \[x \neq -4\] \[x+4=\frac{ 3 }{ y },x=\frac{ 3 }{ y }-4,y \neq 0\]